OpenStudy (anonymous):

simplify radical expression sqrt{?} 80

OpenStudy (radar):

\[\sqrt{80}=\sqrt{5\times16}=\sqrt{16}\sqrt{5}=---------------\]Fill in the blalnk

OpenStudy (radar):

Note that 16 is a perfect square.

OpenStudy (anonymous):

4sqrt{?}5 ?

OpenStudy (radar):

Yes, \[4\sqrt{5}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (radar):

You're welcome